# Spectuner-D1: Spectral Line Fitting of Interstellar Molecules Using Deep Reinforcement Learning

### v1

@@ 2.3.1 The spectral line model
where $J^{bg}_{\nu}$ is the background intensity, $\theta_{\text{maj,min}}$ is the major (minor) axis of the synthesis beam, $c$ is the speed of light and $k$ is the Boltzmann constant. The energy of the lower state $E^{t}_{l}$, the transition frequency $\nu^{t}$, the upper state degeneracy $g^{t}_{u}$, the Einstein coefficient $A^{t}_{\text{ul}}$, and the partition function $Q(T_{\text{ex}})$ are obtained from the spectroscopic database mentioned in Section 2.1.
@@
In addition, we follow the approach in 2017A&A...598A...7M to account for the instrumental resolution effect. The final output model spectrum is obtained by computing the following integral:
@@
2.3.2 The fitting loss
This work adopts the peak-matching loss function developed by 2025ApJS..277...21Q to measure the similarity between the model and observed spectra. This function is also used as the optimization objective for the baseline fitting tasks described in Section 2.7. 2025ApJS..277...21Q demonstrated that the peak-matching loss function can lead to reasonable fitting results in the line-rich regions such as SgrB2(N), significantly more robust than the $\chi^{2}$ function. The peak-matching loss function can be written as
@@ 2.3.2 The fitting loss
where $\ell_{\text{MAE}}$ is the mean absolute error and $\ell_{\text{peak}}$ measures the degree to which the peaks in the model and observed spectra match. The calculation of $\ell_{\text{peak}}$ is complicated, and readers are referred to 2025ApJS..277...21Q for details.
2.4 Training and testing data
@@
The ALMA Three-millimeter Observations of Massive Star-forming regions (ATOMS) project is a band-3 spectral line survey targeting 146 active star-forming regions across the Galactic plane (2020MNRAS.496.2790L). The sample includes hot molecular cores and UC H ii region candidates, with angular resolutions ranging from 1.3′′ to 2.7′′. The data cubes were produced by combining observations of ALMA 7-m array (ACA) and 12-m-array, using CASA version 5.6 (2020MNRAS.496.2790L). The 146 sources are divided into training and testing after preprocessing in Section 2.4.1.
IRAS16293–2422 is a binary system comprising two typical low-mass protostars, IRAS16293 A and IRAS16293 B, located at a distance of approximately 141 pc (2018A&A...614A..20D). The two components exhibit distinct spectral intensities and line widths. When angular resolution permits, they are treated as separate sources for the analysis. We retrieved the data cubes from five archival ALMA projects, processed using the standard pipeline and CASA versions at the time of each observation (2022PASP..134k4501C). The data cover ALMA bands 3, 6, and 7, with angular resolutions ranging from 0.5′′, 1.0′′, and 5.9′′, respectively.
@@
We also include a line survey of Orion-KL obtained from combined Submillimeter Array (SMA) interferometric and IRAM 30 m single-dish observations (2015A&A...581A..71F). It covers 4 GHz at band 6, with a resolution of 4.9′′. 2015A&A...581A..71F converted the 30 m single-dish data into visibilities, and combined them with the SMA data using the MIRIAD package UVMODEL task  [^3] https://www.cfa.harvard.edu/sma/miriad/manuals/SMAuguide/smauserhtml/uvmodel.html.
@@ 2.4.1 Data preprocessing
We correct the local standard of rest (LSR) velocity for cubes to ensure that the velocity offsets of the species are within the fitting range given in Table 1. For the ATOMS data, we adopt the values estimated by 2020MNRAS.496.2790L. For IRAS16293, 2024A&A...686A..59N assumed $v_{\text{LSR}}=2.7$ km/s. We neglect such small LSR velocity and apply no velocity correction to all the IRAS16293 data. In terms of G327, the LSR velocity is roughly between -43.2 to -46.0 km/s (2017A&A...602A..70L; 2020ApJ...904..168B) and we adopt $v_{\text{LSR}}=-45.0$ km/s in this work. 2015A&A...581A..71F corrected the Orion-KL data assuming $v_{\text{LSR}}=7$ km/s, and no further correction is applied in this work.
In addition, continuum data are taken into count, which serve as an input feature for our neural network (see Section 2.1) and are used to compute the model spectrum (see Equation 1).
@@ 2.4.2 Species
Based on the above assumption, this work defines two sets of molecular species, named Mol-1980 and Mol-2010. The former includes species first detected before or in 1980, while the latter includes those detected between 1980 and 2010. We use astromol (2021zndo...5046939M; 2022ApJS..259...30M) to determine the year each molecule was first detected, and list the molecules in Table 3. Isotopologues are included in the corresponding species set regardless of their detection date or whether they have been detected at all. Hyperfine structure lines are excluded. We use only the species in Mol-1980 for training, and those in Mol-2010 to evaluate the generalization ability of our neural network.
@@
Figure 2: Pixel binning for G327-band6-1.9as. Different colors correspond to pixels in different bins. This binning is used for data sampling, as described in Section 2.4.3.
2.4.3 Data sampling
This section explains how the data are sampled during training and producing baseline data in Section 2.7. To generate a data sample, we first randomly select a dataset. The datasets of the ATOMS data contain multiple sources, while other datasets only contain a single source. If an ATOMS dataset is selected, we randomly select a source from it. Next, we randomly choose a spectrum from a pixel in the selected source using the approach described by the subsequent paragraph. Finally, we randomly select one species from the spectroscopic database using the method described in Section 2.4.2.
For most observations, line-rich pixels are concentrated within a compact region. Consequently, if we randomly select a pixel from a line cube, the pixel may not come from a line-rich region and may contain only a few lines. To resolve the issue, we bin the pixels in the line cubes of all spectral windows based on the total number of peaks estimated assuming a prominence threshold of $4\sigma^{\text{global}}_{\text{RMS}}$. The method to find the peaks is described in Section 2.2. For all observations, we use five evenly spaced bins between the minimum and maximum number of peaks. When a spectrum is needed, we first randomly select a bin and then randomly select a pixel within the bin. As an example, Figure 2 illustrates the pixel binning for G327-band7-1.9as.
2.4.4 Data augmentation
Data augmentation is a technique that reduces overfitting and improves generalization by applying random transformations to the training data (e.g. wang2017effectiveness; shorten2019survey). In this study, we apply the following data augmentation strategies:
@@ 2.4.4 Data augmentation
2.5 Network training
We frame the training of our neural network as a single-step reinforcement learning problem. At each training step, the neural network takes the molecular spectroscopic data and the observed spectrum as input, and samples some fitting parameters. A reward value is then assigned by comparing the resulting model spectrum with the observed spectrum. Samples that result in a better match to the observed spectrum receive higher rewards. These reward values serve as training signals for updating the network, a method commonly known as policy gradient (e.g. williams1992simple).
@@ 2.5 Network training
Finally, the third term $\mathcal{L}_{\text{RB}}$ corresponds to the prioritized replay buffer, a technique we introduce to enhance training. A similar idea was presented by horgan2018distributed. The prioritized replay buffer stores the best-fitting samples from previous iterations, which is implemented using a priority queue, a standard data structure in computer science. When adding an item to a full priority queue, the lowest-priority item is automatically removed. In this work, samples with lower $l_{\text{PM}}$ values are assigned higher priority. Therefore, after many iterations, samples with lower $l_{\text{PM}}$ values remain in the priority queue.
We construct the prioritized replay buffer by randomly selecting observed spectra and species for fitting. Each entry in the replay buffer is identified by four attributes: the dataset name, pixel index, target molecule, and augmentation parameters (see Section 2.4.4). These entries maintain a priority queue that stores $l_{\text{PM}}$ values and their associated fitting parameters. All priority queues have a fixed size of 32. For each training dataset except ATOMS-80-band3-2as, we add 16 randomly chosen pixels from each line number bin to the buffer. For ATOMS-80-band3-2as, we instead add 5 pixels per line number bin from each source. These entries are prepared before training. During training, in addition to computing Equation 13, we introduce an extra step to update the prioritized replay buffer, which is found to improve training performance in practice.
@@
Our neural network is trained using 32 V100 GPUs with 256 CPU cores. We adopt a AdamW optimizer (loshchilov2017decoupled) with a weight decay of 0.04 and a learning rate scheduler proposed by (smith2019super). The learning rate first increases from $7.5\times 10^{-7}$ to $7.5\times 10^{-5}$ for 4000 steps and decreases to $7.5\times 10^{-9}$ for 31000 steps according to a cosine function. The training takes 3 - 4 days.
@@
Nelder-Mead Local No gao2012implementing
COBYLA Local No powell1994direct; Zhang_2023
COBYQA Local No rago_thesis
L-BFGS-B Local Yes byrd1995limited
TNC Local Yes nash1984newton
SLSQP Local Yes kraft1988software
PSO Global No wang2018particle
Figure 3: Relation between the mean relative loss and the number of forward modeling runs using different fitting methods. Each dot shows the average metric over all datasets, and the error bars represent the standard deviation over the datasets. The mean relative loss is defined in Section 2.8. The datasets are described in Section 2.4 and summarized in Table 2. Blue dots correspond to results based on traditional optimization methods. Nelder-Mead and SLSQP are classical local optimizers, whereas particle swarm optimization (PSO) is a global optimization algorithm. Red dots show results where a neural network generates the initial guess for the local optimizers. For the methods based on local optimization, the initial guess is given by the best parameters from $N_{\text{init}}=50$ sampled points. The methods to produce these results are introduced in Sections 2.6 and 2.7.
Figure 4: Relation between the mean relative loss and the number of forward modeling runs with different number of initial points generated by the neural network. For each method, from left to right, the dots show the average metric over all datasets with $N_{\text{init}}=10,25,50,100$. The error bars represent the standard deviation over the datasets. The mean relative loss is defined in Section 2.8. The datasets are described in Section 2.4 and summarized in Table 2.
Figure 5: Mean relative loss of neural network-based methods across individual datasets. For each dataset, we present the results from using the neural network alone and from combining the neural network with the SLSQP optimization algorithm. The results are divided into two sets based on the species for fitting, i.e. Mol-1980 and Mol-2010 as described in Section 2.4.2. While Mol-1980 is used for training, Mol-2010 is reserved for testing. The mean relative loss metric is defined in Section 2.8, and the inference method is introduced in Section 2.6.
2.6 Inference methods
@@
This work assesses the performance of six local optimization algorithms from scipy (2020SciPy-NMeth), summarized in Table 4. To prevent parameter divergence, we restrict our selection to algorithms that support bounded parameter ranges. For gradient based optimizers, i.e. L-BFGS-B, TNC, and SLSQP, a three point numerical differentiation formula is used to compute the gradient. We test different $N_{\text{init}}$ and present the results in Section 3. We also evaluate particle swarm optimization (PSO) (e.g. wang2018particle), which can be combined with the neural network by generating all initial particle positions using the outputs of the neural network.
The role of our neural network is to provide an initial guess, and the subsequent fitting process can in principle be applied with any loss function. While the peak matching loss function is employed during training and evaluation, the proposed inference method can also be applied to $\chi^{2}$ fitting. In Section 4, we demonstrate that both the peak matching and $\chi^{2}$ functions can lead to reasonable but different fitting results.
2.7 Producing evaluation data
For evaluation, we compare the proposed neural network-based method with traditional optimization algorithms. First, we construct a sample of observed spectra for fitting using the sampling method described in Section 2.4.3. For each dataset, we randomly select five spectra from each line number bin. The ATOMS data, i.e. ATOMS-80-band3-2as and ATOMS-20-band3-2as, contain multiple sources. Therefore, we use all sources in the datasets, and select only one spectrum per line number bin for each source. We then perform spectral line fitting on the selected observed spectra using different optimization algorithms. The species used for fitting are from both Mol-1980 and Mol-2010, as described in Section 2.4.2. The number of fitting results for each dataset is presented in Table 2.
In this study, three optimization methods are compared with the neural network-based method. First, particle swarm PSO (e.g. wang2018particle) is executed to search the global minimum. The algorithm is run three times to achieve better results. Following 2025ApJS..277...21Q, we terminate the algorithm if there is no improvement over 15 consecutive iterations, with a minimum of 100 iterations. The population size is set to 28. Secondly, we use randomly initialized local optimizers to demonstrate the effectiveness of the neural network. The Nelder-Mead (gao2012implementing) and SLSQP (kraft1988software) algorithms are chosen as representatives of gradient-free and gradient-based methods, respectively. The best result among $N_{\text{init}}$ random samples within the bounds specified in Table 1 is selected as the initial guess for the local optimizers.
2.8 Metrics
@@
where $\ell^{\text{pred}}_{\text{PM}}$ and $\ell^{\text{test}}_{\text{PM}}$ represent the peak matching fitting loss defined in Equation 10. The proposed metric is inspired by the symmetric mean absolute percentage error (SMAPE) (nguyen2019efficient). Since the values of the peak matching function vary substantially across different observed spectra and fitting species, we use a relative metric to ensure comparability of the fitting results. Secondly, the ground truths in the our fitting problems are unknown, and we aim to identify methods that produce lower fitting loss values. Therefore, we avoid absolute-value-based metrics. The proposed metric ranges from -1 to 1, with smaller values indicating better performance. This work adopts the results based on PSO with $N_{\text{trail}}=3$ for $\ell^{\text{test}}_{\text{PM}}$, which in practice produces acceptable fitting loss values.
@@
Figure 6: Comparison of our pixel-level fitting results with xclass. The fitting region is from an ALMA observation of G327, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section 2.4. The fitted molecule is methanol. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section 2.3.1. The first and second rows display our results. The first row uses random initial guesses while the second row employ the neural network to generate initial guesses. The third row illustrates the xclass results. The predictions from our neural network are also used as initial guesses to generate these results. Finally, the fourth row shows the difference between our results and those based on xclass.
@@
Figure 7: Comparison of pixel-level fitting results using the $\chi^{2}$ and peak matching loss functions. The fitting region is from an ALMA observation of G327 at band 6, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section 2.4. The fitted molecule is methanol. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section 2.3.1. The first and second rows display the results using the $\chi^{2}$ and peak matching functions respectively, and the third row shows their difference. The fourth and fifth rows illustrate the fitted spectra for two pixels, with black lines representing the observed spectra. Only a representative frequency range is shown for clarity.
@@
Figure 8: Pixel-level fitting results of CH3OCHO, $v=0$, CH3OCH3, $v=0$, C2H5CN, $v=0$, and C2H3CN, $v=0$. The fitting region is from an ALMA observation of G327 at band 6, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section 2.4. These fitting results are obtained using the SLSQP algorithm, with initial guesses generated using the neural network. The peak matching loss function is adopted in the fittings. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section 2.3.1. For clarity, we only demonstrate results in the central region, and the color scales of each row are different.
3 Results

### v2

@@ II.3.1 The spectral line model
where $J^{bg}_{\nu}$ is the background intensity, $\theta_{\text{maj,min}}$ is the major (minor) axis of the synthesis beam, $c$ is the speed of light and $k$ is the Boltzmann constant. The energy of the lower state $E^{t}_{l}$, the transition frequency $\nu^{t}$, the upper state degeneracy $g^{t}_{u}$, the Einstein coefficient $A^{t}_{\text{ul}}$, and the partition function $Q(T_{\text{ex}})$ are obtained from the spectroscopic database mentioned in Section II.1.
@@
In addition, we follow the approach in Möller et al. (2017) to account for the instrumental resolution effect. The final output model spectrum is obtained by computing the following integral:
@@
II.3.2 The fitting loss
This work adopts the peak-matching loss function developed by Qiu et al. (2025) to measure the similarity between the model and observed spectra. This function is also used as the optimization objective for the baseline fitting tasks described in Section II.7. Qiu et al. (2025) demonstrated that the peak-matching loss function can lead to reasonable fitting results in the line-rich regions such as SgrB2(N), significantly more robust than the $\chi^{2}$ function. The peak-matching loss function can be written as
@@ II.3.2 The fitting loss
where $\ell_{\text{MAE}}$ is the mean absolute error and $\ell_{\text{peak}}$ measures the degree to which the peaks in the model and observed spectra match. The calculation of $\ell_{\text{peak}}$ is complicated, and readers are referred to Qiu et al. (2025) for details.
II.4 Training and testing data
@@
The ALMA Three-millimeter Observations of Massive Star-forming regions (ATOMS) project is a band-3 spectral line survey targeting 146 active star-forming regions across the Galactic plane (Liu et al., 2020). The sample includes hot molecular cores and UC H ii region candidates, with angular resolutions ranging from 1.3′′ to 2.7′′. The data cubes were produced by combining observations of ALMA 7-m array (ACA) and 12-m-array, using CASA version 5.6 (Liu et al., 2020). The 146 sources are divided into training and testing after preprocessing in Section II.4.1.
IRAS16293–2422 is a binary system comprising two typical low-mass protostars, IRAS16293 A and IRAS16293 B, located at a distance of approximately 141 pc (Dzib et al., 2018). The two components exhibit distinct spectral intensities and line widths. When angular resolution permits, they are treated as separate sources for the analysis. We retrieved the data cubes from five archival ALMA projects, processed using the standard pipeline and CASA versions at the time of each observation (CASA Team et al., 2022). The data cover ALMA bands 3, 6, and 7, with angular resolutions ranging from 0.5′′, 1.0′′, and 5.9′′, respectively.
@@
We also include a line survey of Orion-KL obtained from combined Submillimeter Array (SMA) interferometric and IRAM 30 m single-dish observations (Feng et al., 2015). It covers 4 GHz at band 6, with a resolution of 4.9′′. Feng et al. (2015) converted the 30 m single-dish data into visibilities, and combined them with the SMA data using the MIRIAD package UVMODEL task  [^3] https://www.cfa.harvard.edu/sma/miriad/manuals/SMAuguide/smauserhtml/uvmodel.html.
@@ II.4.1 Data preprocessing
We correct the local standard of rest (LSR) velocity for cubes to ensure that the velocity offsets of the species are within the fitting range given in Table 1. For the ATOMS data, we adopt the values estimated by Liu et al. (2020). For IRAS16293, Nazari et al. (2024a) assumed $v_{\text{LSR}}=2.7$ km/s. We neglect such small LSR velocity and apply no velocity correction to all the IRAS16293 data. In terms of G327, the LSR velocity is roughly between -43.2 to -46.0 km/s (Leurini et al., 2017; Beuther et al., 2020) and we adopt $v_{\text{LSR}}=-45.0$ km/s in this work. Feng et al. (2015) corrected the Orion-KL data assuming $v_{\text{LSR}}=7$ km/s, and no further correction is applied in this work.
In addition, continuum data are taken into count, which serve as an input feature for our neural network (see Section II.1) and are used to compute the model spectrum (see Equation 1).
@@ II.4.2 Species
Based on the above assumption, this work defines two sets of molecular species, named Mol-1980 and Mol-2010. The former includes species first detected before or in 1980, while the latter includes those detected between 1980 and 2010. We use astromol (McGuire, 2021, 2022) to determine the year each molecule was first detected, and list the molecules in Table 3. Isotopologues are included in the corresponding species set regardless of their detection date or whether they have been detected at all. Hyperfine structure lines are excluded. We use only the species in Mol-1980 for training, and those in Mol-2010 to evaluate the generalization ability of our neural network.
@@
Figure 2: Pixel binning for G327-band6-1.9as. Different colors correspond to pixels in different bins. This binning is used for data sampling, as described in Section II.4.3.
II.4.3 Data sampling
This section explains how the data are sampled during training and producing baseline data in Section II.7. To generate a data sample, we first randomly select a dataset. The datasets of the ATOMS data contain multiple sources, while other datasets only contain a single source. If an ATOMS dataset is selected, we randomly select a source from it. Next, we randomly choose a spectrum from a pixel in the selected source using the approach described by the subsequent paragraph. Finally, we randomly select one species from the spectroscopic database using the method described in Section II.4.2.
For most observations, line-rich pixels are concentrated within a compact region. Consequently, if we randomly select a pixel from a line cube, the pixel may not come from a line-rich region and may contain only a few lines. To resolve the issue, we bin the pixels in the line cubes of all spectral windows based on the total number of peaks estimated assuming a prominence threshold of $4\sigma^{\text{global}}_{\text{RMS}}$. The method to find the peaks is described in Section II.2. For all observations, we use five evenly spaced bins between the minimum and maximum number of peaks. When a spectrum is needed, we first randomly select a bin and then randomly select a pixel within the bin. As an example, Figure 2 illustrates the pixel binning for G327-band7-1.9as.
II.4.4 Data augmentation
Data augmentation is a technique that reduces overfitting and improves generalization by applying random transformations to the training data (e.g. Wang et al., 2017; Shorten and Khoshgoftaar, 2019). In this study, data augmentation is applied during training using the following strategies:
@@ II.4.4 Data augmentation
II.5 Network training
We frame the training of our neural network as a single-step reinforcement learning problem. At each training step, the neural network takes the molecular spectroscopic data and the observed spectrum as input, and samples some fitting parameters. A reward value is then assigned by comparing the resulting model spectrum with the observed spectrum. Samples that result in a better match to the observed spectrum receive higher rewards. These reward values serve as training signals for updating the network, a method commonly known as policy gradient (e.g. Williams, 1992).
@@ II.5 Network training
Finally, the third term $\mathcal{L}_{\text{RB}}$ corresponds to the prioritized replay buffer, a technique we introduce to enhance training. A similar idea was presented by Horgan et al. (2018). The prioritized replay buffer stores the best-fitting samples from previous iterations, which is implemented using a priority queue, a standard data structure in computer science. When adding an item to a full priority queue, the lowest-priority item is automatically removed. In this work, samples with lower $l_{\text{PM}}$ values are assigned higher priority. Therefore, after many iterations, samples with lower $l_{\text{PM}}$ values remain in the priority queue.
We construct the prioritized replay buffer by randomly selecting observed spectra and species for fitting. Each entry in the replay buffer is identified by four attributes: the dataset name, pixel index, target molecule, and augmentation parameters (see Section II.4.4). These entries maintain a priority queue that stores $l_{\text{PM}}$ values and their associated fitting parameters. All priority queues have a fixed size of 32. For each training dataset except ATOMS-80-band3-2as, we add 16 randomly chosen pixels from each line number bin to the buffer. For ATOMS-80-band3-2as, we instead add 5 pixels per line number bin from each source. These entries are prepared before training. During training, in addition to computing Equation 13, we introduce an extra step to update the prioritized replay buffer, which is found to improve training performance in practice.
@@
Our neural network is trained using 32 V100 GPUs with 256 CPU cores. We adopt a AdamW optimizer (Loshchilov and Hutter, 2017) with a weight decay of 0.04 and a learning rate scheduler proposed by (Smith and Topin, 2019). The learning rate first increases from $7.5\times 10^{-7}$ to $7.5\times 10^{-5}$ for 4000 steps and decreases to $7.5\times 10^{-9}$ for 31000 steps according to a cosine function. The training takes 3 - 4 days.
@@
Nelder-Mead Local No Gao and Han (2012)
COBYLA Local No Powell (1994); Zhang (2023)
COBYQA Local No Ragonneau (2022)
L-BFGS-B Local Yes Byrd et al. (1995)
TNC Local Yes Nash (1984)
SLSQP Local Yes Kraft (1988)
PSO Global No Wang et al. (2018)
Figure 3: Relation between the mean relative loss and the number of forward modeling runs using different fitting methods. Each dot shows the average metric over all datasets, and the error bars represent the standard deviation over the datasets. The mean relative loss is defined in Section II.8. The datasets are described in Section II.4 and summarized in Table 2. Blue dots correspond to results based on traditional optimization methods. Nelder-Mead and SLSQP are classical local optimizers, whereas particle swarm optimization (PSO) is a global optimization algorithm. Red dots show results where a neural network generates the initial guess for the local optimizers. For the methods based on local optimization, the initial guess is given by the best parameters from $N_{\text{init}}=50$ sampled points. The methods to produce these results are introduced in Sections II.6 and II.7.
Figure 4: Relation between the mean relative loss and the number of forward modeling runs with different number of initial points generated by the neural network. For each method, from left to right, the dots show the average metric over all datasets with $N_{\text{init}}=10,25,50,100$. The error bars represent the standard deviation over the datasets. The mean relative loss is defined in Section II.8. The datasets are described in Section II.4 and summarized in Table 2.
Figure 5: Mean relative loss of neural network-based methods across individual datasets. For each dataset, we present the results from using the neural network alone and from combining the neural network with the SLSQP optimization algorithm. The results are divided into two sets based on the species for fitting, i.e. Mol-1980 and Mol-2010 as described in Section II.4.2. While Mol-1980 is used for training, Mol-2010 is reserved for testing. The mean relative loss metric is defined in Section II.8, and the inference method is introduced in Section II.6.
II.6 Inference methods
@@
This work assesses the performance of six local optimization algorithms from scipy (Virtanen et al., 2020), summarized in Table 4. To prevent parameter divergence, we restrict our selection to algorithms that support bounded parameter ranges. For gradient based optimizers, i.e. L-BFGS-B, TNC, and SLSQP, a three point numerical differentiation formula is used to compute the gradient. We test different $N_{\text{init}}$ and present the results in Section III. We also evaluate particle swarm optimization (PSO) (e.g. Wang et al., 2018), which can be combined with the neural network by generating all initial particle positions using the outputs of the neural network.
The role of our neural network is to provide an initial guess, and the subsequent fitting process can in principle be applied with any loss function. While the peak matching loss function is employed during training and evaluation, the proposed inference method can also be applied to $\chi^{2}$ fitting. In Section IV, we demonstrate that both the peak matching and $\chi^{2}$ functions can lead to reasonable but different fitting results.
II.7 Producing evaluation data
For evaluation, we compare the proposed neural network-based method with traditional optimization algorithms. First, we construct a sample of observed spectra for fitting using the sampling method described in Section II.4.3. For each dataset, we randomly select five spectra from each line number bin. The ATOMS data, i.e. ATOMS-80-band3-2as and ATOMS-20-band3-2as, contain multiple sources. Therefore, we use all sources in the datasets, and select only one spectrum per line number bin for each source. We then perform spectral line fitting on the selected observed spectra using different optimization algorithms. The species used for fitting are from both Mol-1980 and Mol-2010, as described in Section II.4.2. The number of fitting results for each dataset is presented in Table 2.
In this study, three optimization methods are compared with the neural network-based method. First, particle swarm PSO (e.g. Wang et al., 2018) is executed to search the global minimum. The algorithm is run three times to achieve better results. Following Qiu et al. (2025), we terminate the algorithm if there is no improvement over 15 consecutive iterations, with a minimum of 100 iterations. The population size is set to 28. Secondly, we use randomly initialized local optimizers to demonstrate the effectiveness of the neural network. The Nelder-Mead (Gao and Han, 2012) and SLSQP (Kraft, 1988) algorithms are chosen as representatives of gradient-free and gradient-based methods, respectively. The best result among $N_{\text{init}}$ random samples within the bounds specified in Table 1 is selected as the initial guess for the local optimizers.
II.8 Metrics
@@
where $\ell^{\text{pred}}_{\text{PM}}$ and $\ell^{\text{test}}_{\text{PM}}$ represent the peak matching fitting loss defined in Equation 10. The proposed metric is inspired by the symmetric mean absolute percentage error (SMAPE) (Nguyen et al., 2019). Since the values of the peak matching function vary substantially across different observed spectra and fitting species, we use a relative metric to ensure comparability of the fitting results. Secondly, the ground truths in the our fitting problems are unknown, and we aim to identify methods that produce lower fitting loss values. Therefore, we avoid absolute-value-based metrics. The proposed metric ranges from -1 to 1, with smaller values indicating better performance. This work adopts the results based on PSO with $N_{\text{trail}}=3$ for $\ell^{\text{test}}_{\text{PM}}$, which in practice produces acceptable fitting loss values.
@@
Figure 6: Comparison of our pixel-level fitting results with xclass. The fitting region is from an ALMA observation of G327, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section II.4. The fitted molecule is methanol. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section II.3.1. The first and second rows display our results. The first row uses random initial guesses while the second row employ the neural network to generate initial guesses. The third row illustrates the xclass results. The predictions from our neural network are also used as initial guesses to generate these results. Finally, the fourth row shows the difference between our results and those based on xclass.
@@
Figure 7: Comparison of pixel-level fitting results using the $\chi^{2}$ and peak matching loss functions. The fitting region is from an ALMA observation of G327 at band 6, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section II.4. The fitted molecule is methanol. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section II.3.1. The first and second rows display the results using the $\chi^{2}$ and peak matching functions respectively, and the third row shows their difference. The fourth and fifth rows illustrate the fitted spectra for two pixels, with black lines representing the observed spectra. Only a representative frequency range is shown for clarity.
@@
Figure 8: Pixel-level fitting results of CH3OCHO, $v=0$, CH3OCH3, $v=0$, C2H5CN, $v=0$, and C2H3CN, $v=0$. The fitting region is from an ALMA observation of G327 at band 6, covering a total bandwidth of $\sim 5$ GHz. The data are described in Section II.4. These fitting results are obtained using the SLSQP algorithm, with initial guesses generated using the neural network. The peak matching loss function is adopted in the fittings. From left to right, the columns show the fitting results for source size, excitation temperature, column density, velocity width and velocity offset. The spectral line model and these fitting parameters are described in Section II.3.1. For clarity, we only demonstrate results in the central region, and the color scales of each row are different.
III Results